Operations Report – February 13th

Crew 221 Operations Report 13Feb2020

SOL: 11

Name of person filing report: Luis D. Monge

Non-nominal systems:
Notes on non-nominal systems:

Generator: on (at 8:20 pm)
Hours run: 12:40
From what time last night: 7:40 pm
To what time this morning: 7:00 am
List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before the generator is run at night) 73%

Diesel Reading – 75%
Station Propane Reading – 48%
Ethanol Free Gasoline – N/A

Water (loft tank): 35 gallons
Water Meter: 148 units
Water (static tank) – 180 gals
Static to Loft Pump used – no
Water in GreenHab – 152.3 gals (water is lower than estimated, please refer to Greenhab report)
Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging: yes

Spirit rover used: no
Hours: (before EVA) : 130.4
Beginning charge: (Before EVA) 100%
Ending charge: (On return from EVA, before recharging)
Currently charging: yes

Opportunity rover used: no
Hours: n/a
Beginning charge: n/a
Ending charge: n/a
Currently charging: yes

Curiosity rover used: yes
Hours: 137.2
Beginning charge: 100%
Ending charge: 78%
Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used
Reason for use: not used
Oil Added? not used
ATV Fuel Used: Gals not used
# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: operational

Summary of Hab operations: nominal.

Summary of GreenHab operations: operational, black corn experiment set to finish tomorrow, now watered with 2l/day. watering of the plants carried out as per the procedure

Summary of ScienceDome operations: one oven is operating in the lab, the incubator is non-functional, microbiology experiments going on and cultures are being grown in agar plates.

Summary of RAM operations: robot assembly completed, radio communication experiment completed.

Summary of any observatory issues: nominal

Summary of health and safety issues: nothing to report.

Questions, concerns, and requests to Mission Support: Spirit was disconnected; EVA crew reconnected it.

Copyright © The Mars Society. All rights reserved. | Main Site